Convert 27 degrees C to F


1
Expert's answer
2022-03-13T15:21:21-0400

Solution.

"0^oC=32^oF;"

"T(^oF)=T(^oC)\sdot1.8+32;"

"T=27^oC;"

"T(^oF)=27\sdot1.8+32=80.6^oF;"

Answer: "27^oC=80.6^oF."